OpenStudy (anonymous):

Find the equation of the line below. Arrange your answer in the form y = mx + b, where b is the constant. The coordinates are (-2,2) and (2,-2)

OpenStudy (igreen):

You can find the slope first. Slope of a line: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) In this case: \(y_2 = -2\) \(y_1 = 2\) \(x_2 = 2\) \(x_1 = -2\) Plug them in: \(m = \dfrac{-2-2}{~~~2+2}\) Add & Subtract: \(m = \dfrac{-4}{~~~4}\) Can you divide that? @madison.bush

OpenStudy (anonymous):

m=-1 @iGreen

OpenStudy (igreen):

Yep, so the slope is -1. In y = mx + b m is the slope, so we can put -1 in: y = -x + b Also, if you draw a line through (-2, 2) and (2, -2), you'll see that they cross through the x and y-axes at (0, 0), which means the y-intercept is 0. So our final equation is: y = -x + 0 or y = -x

OpenStudy (igreen):

Understand? @madison.bush

OpenStudy (anonymous):

yeah. thanks. but my assignment marked it wrong. thats what i put a little while ago and when it was graded it was wrong. @iGreen

OpenStudy (nincompoop):

OR without the aid of a graph and just by pure analysis: after obtaining the slope, \(m = -1\), use the slope intercept form \(y-y_1 = m(x-x_1) \) to solve for y \(y-2 = -1(x-(-2)) \rightarrow y-2 = -x+2 \rightarrow \huge y=-x \)